OpenStudy (anonymous):

solve for x and y: 3x-2y=8 3x+3y= -1

OpenStudy (bahrom7893):

3x=8+2y 8+2y+3y=-1 5y = -1-8 = -9 5y = -9 y = -9/5

OpenStudy (saifoo.khan):

x = 22/15 y = -9/5

OpenStudy (bahrom7893):

3x = 8 - 18/5 = 40/5 - 18/5 = 22/5 x = 22/15